OpenStudy (anonymous):

Prove the statement: For all integers n, if n is odd, then n2 - 3 is even.

OpenStudy (anonymous):

any odd number doubled is even; a even minus a odd is odd. No this statement is NOT TRUE.

terenzreignz (terenzreignz):

I think it's squared, @rrrrr \[n^2 - 3\]

OpenStudy (anonymous):

oh ok then any odd number squared is odd proved here by the units digit: 1^2=1 3^2=9 5^2=25 7^2=49 9^2=81 all of them are odd and you can add tens digit by adding (n0)^2 where n=the tens digit of the number same with the hundreds digit And then a odd minus a odd is even, based on that 1-1=0 mod 2.

terenzreignz (terenzreignz):

@rrrrr I'd have approached it this way... If n is odd, then n = 2k+1, where k is an integer.... \[n^2=(2k+1)^2=4k^2+4k+1=2(2k^2+2k)+1\] Thus proving that \(n^2 \) is also odd. Anyway, good job.